# Comparing Pre-trained Human Language Models: Is it Better with Human Context as Groups, Individual Traits, or Both?

### v2

@@ 4 Experiments
Our research goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined. To this end, we evaluate performances of the models defined in Section 3 on two user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification. We also compare against GPT-2HLC from Soni et al. (2022) as a PLM adapted to the social media domain but devoid of human context.
@@ 4.1 User Level Regression Tasks
We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment Kosinski et al. (2013); Park et al. (2015). These social scientific tasks require predicting continuous outcomes for a user given multiple documents written by them. We use the same data splits as used by Soni et al. (2022) for evaluating HaRT and GPT-2HLC
We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively. We further evaluate these models on the test sets for personality assessment and age estimation, respectively, after continuing training for these tasks, as described in section 3.5.
We use the results from Soni et al. (2022) for HaRT and GPT-2HLC which are directly comparable to GRIT models trained on the same data splits and metrics. Soni et al. fine-tuned the recurrence module of the pre-trained HaRT model for the tasks of age estimation and personality assessment using the average of user-states from non-padded blocks of texts from an author, resulting in two fine-tuned models. Similarly, they fine-tune the last two layers of the pre-trained GPT-2HLC model for these tasks. Since GPT-2HLC can not handle all text from a user in one pass, they average the predictions across all user messages corresponding to the same label for each message.
@@ 4.2 Document-Level Classification Tasks
We compare different models for stance detection vs. topic detection and age classification. These tasks classify a single input document (tweet message or a review) a user writes into label categories. For stance detection, we also use the historical messages of a user where available, as in Soni et al. (2022). We do not have the user information or any user historical language available for the other two tasks, so we evaluate on the single document input.
All models process the input document(s) and feed the layer-normed last non-padded token representation to the classification layer to classify the document into label categories. Only GRIT and HaRT incorporate user information and the historical language (where available). The other two models can only use the input document without a hierarchical structure to make the predictions. We compare with the results from Soni et al. (2022) and Hung et al. (2023) wherever applicable and fine-tune all the parameters of the respective pre-trained models and the classification heads for other task-model combinations using the standard cross-entropy loss.
@@
Given a single annotated tweet, this task predicts a user’s stance as in favor of, against, or neutral towards one of the five targets: atheism, climate change as a real concern, feminism, Hillary Clinton, and legalization of abortion. We fine-tune the models under comparison for each target separately. We report average of weighted F1 scores with three labels across all five targets. We use Soni et al. (2022)’s train/dev/test split over SemEval 2016 dataset (Mohammad et al., 2016). HaRT and GRIT models maintain the temporal accuracy by using only the messages posted earlier than the labeled messages from the extended dataset (Lynn et al., 2019) as a user’s historical language. We compare the results of fine-tuned GPT-2HLC, HaRT (Soni et al., 2022), and fine-tuned GRITage and GRITope.
@@
We use the US subset of the TrustPilot reviews dataset Hovy (2015) from two age groups: below 35 or above 45  [^2] As suggested by Hovy (2015), this split of the age ranges results in roughly equally-sized data sets and is non-contiguous, avoiding fuzzy boundaries.. Given a single review, this task predicts the review topics from five categories: Flights, Online marketplace, Fitness & Nutrition, Electronics, and Hotels. We use the same train, development, and test set splits as Hung et al. (2023) to eliminate any skew in the demographically-conditioned label distribution. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
@@
We use the same subset of the TrustPilot dataset as for topic detection and the same train, development, and test splits from Hung et al. (2023). Given a single review, this task predicts the age group binary label (<35 years old or >45 years old). Age categories are equally represented in each set. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
4.3 Human Language Modeling
5 Results and Discussion
5.1 Comparisons Study
User-Level Regression Tasks.
Table 1 shows the two user-level regression task results. We find that just the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality. Additionally, comparing the transfer learning (Section 3.5) results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find the GRIT models to fare better.
Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it is still devoid of human context. HaRT adds pre-training with the individual human context, and GRIT adds pre-training with both group and individual human contexts (Figure 1). As Table 1 shows, there are gains going from GPT-2HLC (no human context) to HaRT (individual human context) and further to GRIT (individual + group human context). This indicates that pre-training PLMs with individual and group human context may benefit multi-document user-level regression tasks like the ones we considered.
@@ 5.1 Comparisons Study
GPT-2HLC None 68.6 69.8 65.4 63.9
BERTage-MLM Group - 68.4 64.6 $61.9$
BERTDS Group - 69.3 65.0 $64.1$
HaRT Individual $\textbf{71.1}^{*}$ $\textbf{69.8}^{*}$ 65.6 $\textbf{64.3}^{*}$
GRITage Ind + Grp 70.8 69.2 64.5 62.7
GRITope Ind + Grp 70.1 66.5 64.8 61.2
[…]
Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute classification (AC). We see that task fine-tuned HaRT (individual human context) models performs better on all tasks..
HaRT models inherently have an additional context of the individual user and do not treat all inputs as if written by the same user. The tasks considered here relate more to personal opinions and preferences, rather than group-level ones. HaRT model is well-suited for incorporating such personalization due to its pre-training with individual human context. Additionally, the results indicate that the models pre-trained with group context (BERTDS) fare well in the group-based tasks of topic detection and age classification. Whereas, the models pre-trained on both individual and group human context (GRIT) appears to bring in noise for the group-based and personal stance detection tasks resulting in a slightly worse performance.
Further, it is important to note that the individual human context (HaRT) derived from the historical tweets from the users in the stance detection dataset provides for a richer human context. Although, the performance is not greatly hurt even if historical language is not available for certain tasks (TD and AC).
Perplexity.
We also compare the language modeling capability of the various models. Table 3 reports perplexity on the test set of 4438 users. The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT (individual human context) models perform best while GRIT (individual + group human context) models result in a slightly lower perplexity than HaRT. GRIT models are pre-trained under a multi-task learning setup which most likely is hurting the individual task performances, thus resulting in a slight dip in the perplexity compared to HaRT. Further, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al. (2022).
@@
5.2 Error Analysis and Disparity
We conduct an error analysis as a function of a socio-demographic group attribute. Specifically, we measure GRIT and HaRT for error disparity (Shah et al., 2020) – – a systematic difference in error as a function of demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015). We analyzed outcome and error disparity on age and openness prediction from both the models with individual context (HaRT) and combined individual plus group context (GRIT).
First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set. Then, we compare the performance of our models across these buckets. Results from table 4 indicate that pre-training with individual and group context together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age. We see similar trends for assessing openness personality (Appendix Tables 6, 8). This suggests that the group attribute prediction maybe acting as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, that is averaged by the number of pairs (Shah et al., 2020). A lower averaged sum of differences implies lesser errors as a function of the age groups. Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
@@ 5.2 Error Analysis and Disparity
6 Related Work
Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models. Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups. These works improved sentiment analysis, polarity classification, and topic detection. Other researchers explored adapting to user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) to improve downstream tasks like sarcasm detection, and toxic language detection. Some studies go beyond explicit groups and learning individual representations latently Jaech and Ostendorf (2018); Delasalles et al. (2019) or via historical language Matero et al. (2021).
With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks. More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context. However, few studies have explored including human context within the pre-training regime of LMs. Hung et al. (2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup. Further, Soni et al. (2022) pre-train a LM with individual human context derived from user language. Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
@@ 7 Conclusion
NLP benefits from modeling latent human context, such as sociodemographic group traits or individual tendencies. A recent development has been to incorporate this additional human context into the PLMs’ pre-training regimen. However, humans exhibit varying degrees of group and individual characteristics. Learning about the impacts of pre-training with different types of human context will propel us forward in including human context in our base LLMs (Soni et al., 2023). To assess its impacts, we compare three types of PLMs pre-trained with sociodemographic group attributes, individual human contexts, or a combination across five user- and document-level tasks. We find that pre-training with individual and group human context improves the two user-level tasks: age and personality prediction. Pre-training with individual human context improves the three single-document classification tasks, including stance and topic detection. Pre-training solely on group context helps group-based document classification tasks, though not optimally. Our findings show a promising step towards modeling human context and provide valuable insights for the NLP community to think of additional strategies for improving models with task-dependent human context in pre-training.
@@
The purpose of our study is to compare the impacts of modeling sociodemographic group attributes and modeling individual user traits, and we use relevant models to represent each of the approaches. There are likely to be other ways to model these approaches and the models we use are only one of the ways. Additionally, these models in themselves have limitations like the blocks mechanism to process all the text from author induces compute requirements resulting in a capping of the number of blocks used for training. While it is also unclear how many blocks are sufficient to capture the human context, and if it is helpful to use the earliest language or the most recently used language in the capped number of blocks.
@@
Models that incorporate sociodemographic information need to be considered with special scrutiny. On the one hand, they have the potential to produce fairer and more inclusive results, because they can account for human language variation. On the other hand, they risk revealing identifying or sensitive information, which can lead to profiling and stereotyping. These may present opportunities for unintended malicious exploitations. For example, models that improve demographic groups prediction or psychological assessments could be used for targeting content for individuals without their awareness or consent. Such models may also risk release of private information of the research participant if trained on private data unchecked for exposing identifying information. For this reason, we take a conservative release strategy. While we support open research and reproducibility, data and privacy protection take precedence. Thus, we will only be releasing the code for our comparison study and the data that does not contain sensitive information i.e., stance detection datasets and TrustPilot datasets for topic detection and attribute classification. This is also in accordance with the DUA we have received from the authors of the papers/models that we employ in our work.
@@ Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #585968. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.

### v3

@@ 5 Experiments
Our study’s goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined. To this end, we evaluate performances of the models defined in Section 4 on two multi-document user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification. We also compare against GPT-2HLC from Soni et al. (2022) as a PLM adapted to the social media domain but devoid of human context. All experiments were run using Optuna trials (Akiba et al., 2019) to search for the best hyperparameters and reduce the effects of randomness. More details are provided in Appendix A.2
@@ 5.1 User Level Regression Tasks
We consider two user-level social scientific tasks: age estimation, and personality (openness) assessment, which require predicting continuous outcomes (real-valued age, or openness score) for a user given multiple documents written by them. We use the same data splits as used by Soni et al. (2022) for our comparison study.
Since GRITage is pre-trained using age estimation as one of the tasks, we use directly evaluate it on the held-out test set. This allows for direct comparison with HaRT fine-tuned for the age estimation task. Furthermore, we can potentially attribute performance differences to the training with combined group and individual context, as GRITage incorporates the group feature into HaRT’s architecture. Similarly, GRITope is evaluated on the held-out test set for personality assessment. Moreover, we evaluate GRITage and GRITope for the tasks of personality assessment and age estimation, respectively, using the transfer learning mechanism described in Section 4.5. We report and compare pearson $r$ for age estimation and disattenuated pearson $r$ for personality assessment.
@@ 5.2 Document-Level Classification Tasks
We compare different models for stance detection vs. topic detection and age classification tasks. These tasks classify a single input document (tweet message or a review) that a user writes into label categories. For stance detection, we also utilize the historical messages of a user where available, as in Soni et al. (2022). However, we do not have the user information or any user historical language available for the other two tasks, so we evaluate solely based on the single document input.
All models process the input document(s) and feed the layer-normed last non-padded token representation to the classification layer to classify the document into label categories. Only GRIT and HaRT incorporate user information and the historical language available for the stance detection task. However, GPT-2HLC, and both BERTDS and BERTage-MLM lack this hierarchical structure and can only use the input document without access to historical data for making predictions. We compare the results from Soni et al. (2022) and Hung et al. (2023) wherever applicable and fine-tune all the parameters of the respective pre-trained models and the classification heads for other task-model combinations using the standard cross-entropy loss.
@@
Given a single annotated tweet, this task predicts a user’s stance as in favor of, against, or neutral towards one of the five targets: atheism, climate change as a real concern, feminism, Hillary Clinton, and legalization of abortion. We fine-tune GRITage and GRITope for each target separately, and use the results from Soni et al. (2022) for GPT-2HLC and HaRT. We report the average of weighted F1 scores [^3] We use this metric to maintain consistency with previous works under comparison (Soni et al., 2022; Hung et al., 2023). with three labels across all five targets. We use the train/dev/test split provided by Soni et al. (2022) over the SemEval 2016 dataset (Mohammad et al., 2016). HaRT and GRIT models maintain the temporal accuracy by using only the messages posted earlier than the labeled messages from the extended dataset (Lynn et al., 2019) as a user’s historical language.
@@
We use the US subset of the TrustPilot reviews dataset Hovy (2015) from two age groups: below 35 or above 45  [^4] As suggested by Hovy (2015), this split of the age ranges results in roughly equally-sized data sets and is non-contiguous, avoiding fuzzy boundaries.. Given a single review, the task is to predict the review topics from five categories: Flights, Online marketplace, Fitness & Nutrition, Electronics, and Hotels. To maintain consistency, we adopt the same train, development, and test set splits as Hung et al. (2023) to ensure a stratified demographically-conditioned label distribution. We fine-tune GPT-2HLC, HaRT, GRITage, and GRITope using these data splits to predict the topic for a given review, and report macro-F1 scoresfootnotemark: . We also compare to results from BERTage-MLM and BERTDS (Hung et al., 2023).
@@
We use the same subset of the TrustPilot dataset as for topic detection and the same train, development, and test splits from Hung et al. (2023). Given a single review, this task predicts the age group binary label (<35 years old or >45 years old). Age categories are equally represented in each set. We fine-tune GPT-2HLC, HaRT, GRITage and GRITope using the provided splits to predict if the review is written by someone below 35 years or above 45 years, and report macro-F1 scoresfootnotemark: . We also compare to results from BERTage-MLM and BERTDS (Hung et al., 2023).
@@
5.3 Human Language Modeling
GPT-2HLC None 68.60 69.77 65.43 63.93
BERTage-MLM Group - 68.40 64.60 61.90
BERTDS Group - 69.30 65.00 64.10
HaRT Individual 71.10 69.84 65.65 64.33
GRITage Ind + Grp 70.82 69.21 64.52 62.56
GRITope Ind + Grp 70.07 66.53 64.84 61.18
Table 2: Weighted F1 for stance detection, macro-F1 for topic detection (TD), and age classification (AC) on TrustPilot reviews. Pre-training with individual context appear to benefit all tasks. Bold $=$ best in column; McNemar’s test comparing classifiers does not show statistical significance between the best performing model (HaRT) and the best baseline with no individual context (GPT-2HLC).
6 Results and Discussion
6.1 Comparisons Study
User-Level Regression Tasks.
Table 1 shows the results of the two user-level regression tasks. We find that GRIT models outperform others for both age estimation and personality assessment tasks. Additionally, upon comparing the transfer learning (Section 4.5) outcomes of GRITage for openness and GRITope for age to those of the HaRT and GPT-2HLC models, we consistently observe superior performance with the GRIT models, further substantiating their efficacy.
Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it lacks human context. HaRT incorporates individual human context in pre-training, and GRIT extends this by integrating both group and individual human contexts in pre-training (Figure 1). As shown in Table 1, there are gains observed from GPT-2HLC (no human context) to HaRT (individual human context), and further to GRIT (individual + group human context). This suggests that pre-training PLMs with individual and group human context can benefit multi-document user-level regression tasks, such as those we considered. Importantly, the only difference between HaRT and GRIT models lies in the integration of the demographic attribute prediction (group context). Both models are pre-trained and evaluated on precisely the same data, allowing performance differences to be attributed to the additional group context combined with individualistic human context.
@@ 6.1 Comparisons Study
Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute (age) group classification (AC). We see that task fine-tuned HaRT (individual human context) models perform better on all tasks.
HaRT models inherently include an additional context of the individual user and do not treat all inputs as if written by the same user. The considered stance detection task primarily relates to personal opinions and preferences, rather than group-level ones, making HaRT well-suited for incorporating such personalization due to its pre-training with individual human context. While a group context may also influence a person’s stance to some extent, empirical observations show that the combination of individual and group contexts negatively impacts performance. Additionally, models pre-trained with group context (BERTDS) perform well in group-based tasks such as topic detection and age classification. However, models pre-trained on both individual and group human context (GRIT) do not appear to enhance results in group-based, and personal stance detection tasks resulting in slightly worse performance.
Further, it is important to note that the individual human context (HaRT) derived for some of the users using their historical tweets, where available, in the stance detection dataset provides a richer human context as we see greater gains in the performance of HaRT over GPT-2HLC. Conversely, when historical language is not available for certain datasets (topic detection and attribute classification), HaRT does not perform worse than GPT-2HLC and may even achieve marginal gains due to the inherent human context in the model. However, we leave the evaluation of the impact of historical language on human context for future work.
@@
Perplexity.
We also compare the language modeling capability of the various models. Table 3 reports perplexity on the held-out test set. The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, while HaRT model with individual human context perform the best. In contrast, GRIT models with both individual and group human context exhibit a slightly lower perplexity than HaRT. An individual’s language is inherently personal, yet it can also be influenced by their group context to some extent, thereby affecting the perplexity results in language modeling tasks. However, GRIT models pre-trained with both individual and group context yield slightly worse perplexity measures. Additionally, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al. (2022).
@@
6.2 Error Analysis and Disparity
We conduct an error analysis based on a socio-demographic group attribute (age groups), specifically focusing on age and openness prediction tasks. We measure the performance of GRIT and HaRT in terms of error disparity (Shah et al., 2020) — a systematic difference in error based on demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015). We analyze both the prediction outcomes and error disparity in age and openness prediction for both models: HaRT, which considers individual context, and GRIT, which incorporates both individual and group context.
First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set, and then we compare the performance of our models across these buckets. Results from Table 4 indicate that pre-training with individual and group contexts together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age. We see similar trends for assessing openness personality (see Appendix Tables 6 and 8), suggesting that the group attribute prediction may act as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, which is then averaged by the number of pairs (Shah et al., 2020). A lower averaged sum of differences implies fewer errors as a function of the age groups. Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
@@ 7 Conclusion
NLP benefits from modeling latent human context, such as socio-demographic group features or individual traits. A recent development has been to incorporate this additional human context into the pre-training regimen of LMs. However, humans exhibit varying degrees of group and individual characteristics. Understanding the impacts of pre-training with different types of human context will advance the integration of human context into our base LLMs (Soni et al., 2024). To assess the impacts, we compare three types of PLMs pre-trained with socio-demographic group attributes, individual human contexts, and combined group and individual traits, across five user- and document-level tasks. Our findings indicate that pre-training with both individual and group human context improves the two user-level regression tasks: age and personality prediction. Pre-training with individual human context enhances the performance of the three single-document classification tasks, including stance and topic detection. Interestingly, inclusion of both individual and group attributes results in reduced performance on the text classification tasks. Meanwhile, pre-training solely on group context aids in group-based document classification tasks, albeit suboptimally. These results represent a promising step towards modeling human context and offer valuable insights for the NLP community to investigate additional strategies for improving models with task-dependent human context during pre-training.
@@
The purpose of our study is to compare the impacts of modeling socio-demographic group attributes and modeling individual user traits, and we use relevant models to represent each of the approaches. There are likely to be other ways to model these approaches and the models we use are only one of the ways. Additionally, these models in themselves have limitations like the blocks mechanism to process all the text from author induces compute requirements resulting in a capping of the number of blocks used for training. While it is also unclear how many blocks are sufficient to capture the human context, and if it is helpful to use the earliest language or the most recently used language in the capped number of blocks.
@@
Models that incorporate socio-demographic information need to be considered with special scrutiny. On the one hand, they have the potential to produce fairer and more inclusive results, because they can account for human language variation. On the other hand, they risk revealing identifying or sensitive information, which can lead to profiling and stereotyping. These may present opportunities for unintended malicious exploitations. For example, models that improve demographic groups prediction or psychological assessments could be used for targeting content for individuals without their awareness or consent. Such models may also risk release of private information of the research participant if trained on private data unchecked for exposing identifying information. For this reason, we take a conservative release strategy. While we support open research and reproducibility, data and privacy protection take precedence. Thus, we will only be releasing the code for our comparison study and the data that does not contain sensitive information i.e., stance detection datasets and TrustPilot datasets for topic detection and attribute classification. This is also in accordance with the DUA we have received from the authors of the papers/models that we employ in our work.
@@ Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005, and a grant from the CDC/NIOSH (U01 OH012476). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, any other government organization, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.